OpenStudy (anonymous):

area between y^2-2x=5 and x-y=5

OpenStudy (anonymous):

my first step was to get them both in terms of y so I have \[\frac{ 1 }{ 2 }y ^{2}-\frac{ 5 }{ 2 }\]

OpenStudy (anonymous):

and y+5

OpenStudy (anonymous):

I set them equal and found the intersection between the two to be -3 and 5

OpenStudy (anonymous):

I integrated \[\int\limits_{-3}^{5}(y+5)-(\frac{ 1 }{ 2 }y ^{2}-\frac{ 5 }{ 2 })dy\]

OpenStudy (anonymous):

and I keep getting 128/3 which is not any of my answer choices

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@dan815

OpenStudy (phi):

Here is a plot of the figure. As you can see some of the area is below the x-axis and if you compute "blindly" will be treated as *negative* i.e. subtracted from the top area.

OpenStudy (phi):

Good news/Bad news: I got 128/3

OpenStudy (phi):

Your answer is correct, so you might want to talk to your teacher about this one. out of curiosity, what are the answer choices?

OpenStudy (anonymous):

I no longer have access to the answer choices but it says the correct answer is suppose to be 196/3